Virginia bluebells have flowers that can be blue or white. The data below contains data on pollination rates for plants (numpoll) with either blue or white flowers (color), and the number of flowers on each plant (plsize). Do the regressions of number of flowers pollinated on plant size (number of flowers) for blue and white flowers have the same slopes and intercepts? Are the regressions coincident or parallel? PLS USE R SYNTAX/R MARKDOWN

color plsize numpoll colorcode
white 6 2 0
white 12 9 0
white 22 11 0
white 14 11 0
white 16 9 0
white 13 10 0
white 11 8 0
white 8 4 0
white 4 1 0
white 24 14 0
white 5 4 0
white 15 11 0
blue 24 19 1
blue 4 4 1
blue 9 5 1
blue 14 12 1
blue 17 13 1
blue 12 10 1
blue 10 7 1
blue 6 5 1
blue 18 15 1
Answer #1

R codes:

> w_plsize=scan('clipboard');w_plsize
Read 12 items
[1] 6 12 22 14 16 13 11 8 4 24 5 15
> w_numpoll=scan('clipboard');w_numpoll
Read 12 items
[1] 2 9 11 11 9 10 8 4 1 14 4 11
> b_plsize=scan('clipboard');b_plsize
Read 9 items
[1] 24 4 9 14 17 12 10 6 18
> b_numpoll=scan('clipboard');b_numpoll
Read 9 items
[1] 19 4 5 12 13 10 7 5 15
> model1=lm(w_numpoll~w_plsize)
> summary(model1)

Call:
lm(formula = w_numpoll ~ w_plsize)

Residuals:
Min 1Q Median 3Q Max
-2.48551 -1.31102 -0.02326 1.51793 2.27422

Coefficients:
Estimate Std. Error t value Pr(>|t|)
(Intercept) 0.39626 1.16862 0.339 0.742
w_plsize 0.59497 0.08419 7.067 3.43e-05 ***
---
Signif. codes: 0 ‘***’ 0.001 ‘**’ 0.01 ‘*’ 0.05 ‘.’ 0.1 ‘ ’ 1

Residual standard error: 1.76 on 10 degrees of freedom
Multiple R-squared: 0.8332, Adjusted R-squared: 0.8165
F-statistic: 49.94 on 1 and 10 DF, p-value: 3.429e-05

> model2=lm(b_numpoll~b_plsize)
> summary(model2)

Call:
lm(formula = b_numpoll ~ b_plsize)

Residuals:
Min 1Q Median 3Q Max
-2.0482 -0.4885 0.3669 0.7065 0.9769

Coefficients:
Estimate Std. Error t value Pr(>|t|)
(Intercept) -0.19706 0.83418 -0.236 0.82
b_plsize 0.80503 0.05962 13.503 2.87e-06 ***
---
Signif. codes: 0 ‘***’ 0.001 ‘**’ 0.01 ‘*’ 0.05 ‘.’ 0.1 ‘ ’ 1

Residual standard error: 1.063 on 7 degrees of freedom
Multiple R-squared: 0.963, Adjusted R-squared: 0.9577
F-statistic: 182.3 on 1 and 7 DF, p-value: 2.87e-06

> plsize=c(w_plsize,b_plsize)
> numpoll=c(w_numpoll,b_numpoll)
> plot(plsize,numpoll)
> abline(model1)
> abline(model2)

15 numpoll 10 ooooo O 5 B 10 15 20 plsize

Slope and intercept of the two models namely for white and blue flowers are different.

Regression line do not coincide with each other.